OpenStudy (anonymous):

Five times the difference of a number and 17 is at most 115. Let x represent the number and find all possible values for the number.

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

Hi there. Want to do this out step by step?

OpenStudy (anonymous):

sure...that would be great!

OpenStudy (anonymous):

Alright, so from the beginning: the difference of a number and 17 is what?

OpenStudy (anonymous):

x - 17

OpenStudy (anonymous):

I'll be back in a little bit, sorry! Try and see what you can figure out yourself.

OpenStudy (anonymous):

Can someone help me??

Parth (parthkohli):

Sure :)

OpenStudy (anonymous):

Hey, stil there?

OpenStudy (anonymous):

HI thank you!!!!!!!!!!!!

OpenStudy (anonymous):

yeah

Parth (parthkohli):

If a number is “at most 115”, how would you represent the inequality associated?

OpenStudy (anonymous):

Well, I'll let them take over

OpenStudy (anonymous):

k

OpenStudy (anonymous):

less than or equal to 115

Parth (parthkohli):

Exactly, so how would you represent the whole equality? =)

Parth (parthkohli):

Thank you @srossd.

OpenStudy (anonymous):

x <= 115

Parth (parthkohli):

Well, not really. You have the difference of \(x\) and \(17\) less than or equal to \(115\).

OpenStudy (anonymous):

k

Parth (parthkohli):

So what did you get?

OpenStudy (anonymous):

x < 6

Parth (parthkohli):

Yes :)

Parth (parthkohli):

\[x \le 6\]

OpenStudy (anonymous):

oh thx... @ayushprasad

OpenStudy (anonymous):

thanks for the tip!!!!

OpenStudy (anonymous):

wait

Parth (parthkohli):

Oh no.

OpenStudy (anonymous):

so tell me ... how did u get the answere???

OpenStudy (anonymous):

hi @srossd

Parth (parthkohli):

\[5(x - 17) \le 115\]\[x - 17 \le 23\]Sorry...

OpenStudy (anonymous):

Hi, just kept getting the little alerts for this post and got curious. I'll leave now :).